27. The boy with mass m1 = 52kg is located on the cart with mass m2 = 20kg and from the resting position he shoots horizontally. The mass of lead is m0 = 120g, while v0 = 800m / s. Determine the speed that the cart has taken after five shots. Friction forces are absent.


1
Expert's answer
2021-02-04T02:33:39-0500

According to the law of conseration of linear momentum, we get:


"p_i=p_f,""(m_1+m_2)v_{cart}+5m_0v_0=0,""(m_1+m_2)v_{cart}=-5m_0v_0""v_{cart}=\dfrac{-5m_0v_0}{(m_1+m_2)},""v_{cart}=\dfrac{-5\cdot0.12\ kg\cdot800\ \dfrac{m}{s}}{(52\ kg+20\ kg)}=-6.7\ \dfrac{m}{s}."

The sign minus means that the cart moves in the opposite direction to the motion of the bullets.